Pottery Ranch Inc. has been manufacturing its own finials for its curtain rods. The company is currently operating at 100% of capacity, and variable manufacturing overhead is charged to production at the rate of 54% of direct labor cost. The direct materials and direct labor cost per unit to make a pair of finials are $3.61 and $4.85, respectively. Normal production is 25,100 curtain rods per year.

A supplier offers to make a pair of finials at a price of $13.43 per unit. If Pottery Ranch accepts the supplier’s offer, all variable manufacturing costs will be eliminated, but the $45,300 of fixed manufacturing overhead currently being charged to the finials will have to be absorbed by other products.

(a)

Prepare an incremental analysis to decide if Pottery Ranch should buy the finials. (Round answers to 0 decimal places, e.g. 1250. Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Make Buy Net Income
Increase (Decrease)
Direct materials $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places
Direct labor enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places
Variable overhead costs enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places
Fixed manufacturing costs enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places
Purchase price enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places
Total annual cost $enter a total amount for this column rounded to 0 decimal places $enter a total amount for this column rounded to 0 decimal places $enter a total amount for this column rounded to 0 decimal places


(b)

Should Pottery Ranch buy the finials?

select an option                                                          NoYes, Pottery Ranch should select an option                                                          buynot buy the finials.


(c)

Would your answer be different in (b) if the productive capacity released by not making the finials could be used to produce income of $67,010?

select an option                                                          YesNo, income would select an option                                                          decreaseincrease by $enter a dollar amount
Answer #1
a
Make Buy Net income Increase
(Decrease)
Direct materials 90611 90611
Direct labor 121735 121735
Variable overhead costs 65737 65737
Fixed manufacturing costs 45300 45300 0
Purchase price 0 337093 -337093
Total annual costs 323383 382393 -59010
b
No, Pottery Ranch should not buy the finials
c
Change in income = 67010-59010 = $8000
Yes income would increase by $8000
Formulas used:
Make Buy
Direct materials =25100*3.61
Direct labor =25100*4.85
Variable overhead costs =121735*54%
Purchase price =25100*13.43